OpenStudy (anonymous):

Simplify the expression 4x(2x2 − 7x + 3).

OpenStudy (ivettef365):

first multiply 4x(2x^2) = 8x^3 4x(-7x) = -28x^2 4x(3) = 12x

OpenStudy (ivettef365):

Therefore you have 8x^3-28x^2+12x

OpenStudy (anonymous):

thanks sugar

OpenStudy (anonymous):

srry if your a dude im from texas